# Explicit Min-wise Hash Families with Optimal Size

### v1

@@ 2 Preliminaries
Explicit constructions of $t$-wise independent family of seed length $O(t\log NM)$ are well known. We state two useful bounds on $t$-wise independent random variables [Ind01, FPS11]. In the rest of this work, we use $\operatorname*{\mathbb{E}}_{X:t\text{-wise}}[f(X)]$ denote the expectation of $f(X)$ when $X$ is $t$-wise independent.
@@
PRGs for combinatorial rectangles have been extensively studied in the past few decades [ASWZ96, LLSZ97, Lu02, GMR+12, GY20]. Our constructicons of min-wise hash are based on these PRGs and their techniques. We state the state-of-the-art here by Gopolan and Yehudayoff, i.e., Theorem 1.9 in [GY20].
@@ 3 Min-wise Hash of Polynomial Size
The term in (3.3) is the error introduced by $\mathsf{PRG}_{1}$, which is multiplicative with respect to $\Pr[\sigma(y)=\theta]$. So the last piece is to show the total error of (3.3) over $\theta$, $\sum_{\theta\in[M]}\Pr_{\sigma\sim\mathcal{G}}[\sigma(y)=\theta\wedge\min\sigma(B_{j})>\theta]\cdot 2^{-C_{s}\cdot t}$, is bounded by $o(\delta)/|X|$. The observation here is that $\sum_{\theta\in[M]}\Pr_{\sigma\sim\mathcal{G}}[\sigma(y)=\theta\wedge\min\sigma(B_{j})>\theta]$ is the exact probability of event $(\sigma(y)<\min\sigma(B_{j}))$ under $(0.1C_{s}+1)$-wise independence since $\mathcal{G}$ is $(0.1C_{s}+1)$-wise independent. Thus by the classical result of Indyk [Ind01], this part is bounded by $O(1)/|B_{j}|$.
@@
Theorem 3.5 (Theorem 1.1 in [Ind01]).
@@
    After a union bound, with probability $1-1/\ell^{3C}$, $|B_{i}|\leq\ell^{0.1}+\operatorname*{\mathbb{E}}[|B_{i}|]=2\ell^{0,1}$ for all $i$.
@@ 3.3.1 The First Part
We show (3.5) via the Nisan-Zuckerman framework. One subtle thing is that the extractor error $\mathsf{ExtErr}=2^{-C_{s}\cdot t}$ in (3.5) is multiplicative with respect to $\Pr_{\sigma:\text{$(0.1C_{s}+1)$-wise}}[\sigma(y)=\theta]=1/M$, different from applying the PRG fooling combinatorial rectangles directly. This is shown by permuting the order of $s_{1},\ldots,s_{\ell}$ such that $z_{j}:=\mathsf{Ext}(w,s_{j})$ is the first input of the read-once branching programming.
Given the allocation $g$, we consider a width-$2$ length-$\ell$ read-once branching program $P$ whose input alphabet is $\{0,1\}^{0.3C_{e}\cdot\log N}$ corresponding to $z_{1},\ldots,z_{\ell}$. Because of the definition $B_{i}:=\{x:g(x)=i\}$ and $h^{\prime}(i):=\sigma_{g(i)}(i)$ for $\sigma_{1}:=G_{z_{1}},\ldots,\sigma_{\ell}=G_{z_{\ell}}$, we fix $g$ and rewrite $(h^{\prime}(y)=\theta\wedge\min h^{\prime}(X\setminus y)>\theta)$ as the conjunction of $\ell$ events depending on $z_{1},\ldots,z_{\ell}$:
@@ 3.3.2 The Second Part
$(1-\theta/M)^{|B_{i}|}\pm\varepsilon=(1-\theta/M)^{|B_{i}|}\cdot(1\pm 2^{-(0.5C_{s}-1)\cdot t}),\ \forall i\in[\ell].$
@@
$\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)=(1\pm 2^{-(0.5C_{s}-2)\cdot t})\cdot\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}(1-\theta/M)^{|X|-1}.$
@@
$\dfrac{1}{M}\sum_{(1-\theta/M)^{C_{g}}\leq 2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)\leq 2^{-0.5C_{s}\cdot t/C_{g}}.$
@@
$=$ $\dfrac{1}{M}\left(\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)+\sum_{(1-\theta/M)^{C_{g}}\leq 2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)\right)\pm 1/\ell^{3C}$
$=$ $\frac{1\pm 2^{-(0.5C_{s}-2)\cdot t}}{M}\cdot\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}(1-\theta/M)^{|X|-1}\pm 2\cdot 2^{-0.5C_{s}\cdot t/C_{g}}\pm 1/\ell^{3C}.$
@@
$2^{-(0.5C_{s}-2)\cdot t}+\dfrac{1}{\ell^{3C-0.9}}+2|X|\cdot 2^{-0.5C_{s}\cdot t/C_{g}}\leq\dfrac{1}{\ell^{3C-0.9}}+2^{-0.5C_{s}\cdot t}+2^{-(0.5C_{s}/C_{g}-1)\cdot t}<2^{-2C\cdot t}.$
@@
$=$ $\dfrac{1}{M}\sum_{\theta\leq 0.5M\cdot\ell^{-0.2}}(1-\theta/M)^{|X|-1}\cdot\left(1\pm 2^{-0.5C_{s}\cdot t}\right).$
@@
    $\leq$ $2^{2S-C^{\prime}\cdot(C_{s}-1)\cdot t\cdot S}=N^{-O(1)},$
    where we plug the definition of $S^{\prime}=C^{\prime}\cdot\log\log N$ in the last step.
@@
$\dfrac{1}{\ell^{3C-1.1}}+2^{-0.5C_{s}\cdot t}+2|X|\cdot\left(\exp(-\ell^{0.6})+N^{-O(1)}\right)<2^{-2C\cdot t}.$
@@
where the last line applies the fact that only those $s_{j}=\alpha$ which make $(\sigma_{j}(\theta)=\theta\wedge\min\sigma_{j}(B_{j})>\theta)$ true have contributions to the expectation.
@@ 4 $k$-min-wise Hash
We use the second approach outlined in Section˜1.2 to construct $k$-min-wise hash in this section. Recall the definition that $\mathcal{H}=\{h_{i}:[N]\to[M]\}$ is a $k$-min-wise hash family of error $\delta$, if for any $X\subseteq[N]$ and $Y\subseteq X$ of size at most $k$, $\Pr_{h\sim\mathcal{H}}[h(y)<\min h(X\setminus y)]=\frac{1\pm\delta}{{|X|\choose|Y|}}$. Without loss of generality, we assume $|Y|=k$ in this section.
@@
If we neglect the $2^{-C_{s}\cdot t\cdot k+1}$ factor and consider $\sum_{\theta\in[M]}\Pr_{\sigma:(C_{e}+1)\cdot k\text{-wise}}[\max\sigma(Y)=\theta\penalty 10000\ \wedge\penalty 10000\ \min\sigma(B_{J})>\theta]$, this is the exact probability of $B_{J}\cup Y$ satisfying the $k$-min-wise hash condition under $(C_{e}+1)\cdot k$-wise independence. Feigenblat, Porat and Shiftan [FPS11] have shown this bounded by $2k!/|B_{J}|^{k}$ when $C_{e}$ is large enough. We state their results for completeness.
Theorem 4.6 (Theorem 1.1 in [FPS11]).
@@ 5 Extractors
We show how to design explicit extractors for Lemma˜5.1 in the rest of this section. Our construction will be based on the linear form of the lossless condenser in [GUV09, CI17].
@@
We strengthen the linear seeded condensers in [GUV09, CI17] such that they are surjective for every seed. We state the main properties of lossless condensers as follows, which reformulates Corollary 38 of [CI17] with an extra surjective guarantee.
@@
Let $\alpha>0$ be any constant. For any $n$, any $k<n$, and $\varepsilon>0$, there exists a $(k,\varepsilon)$-lossless condenser $h:\mathbf{F}_{2}^{n}\times\mathbf{F}_{2}^{d}\to\mathbf{F}_{2}^{m}$ with $d\leq(1+1/\alpha)\cdot\log(nk/\varepsilon)+O(1)$ and $m\leq d+(1+\alpha)k$. Moreover, this lossless condenser satisfies the following two properties:
@@
Corollary 38 in [CI17] provides a linear lossless condenser with the above parameters although it is not necessarily surjective. In particular, its construction provides a condenser for every $n$ and every $k<n$ as long as there exists an irreducible univariate polynomial of degree $n$ in some finite extension field of $\mathbf{F}_{2}$. Irreducible polynomials of degree $n$ always exist because the number of irreducible polynomials of degree $n$ in the finite field $\mathbf{F}_{q}$ is $\frac{1}{n}\sum_{d\mid n}\mu(n/d)q^{d}$ greater than 0 (by the Gauss formula). Moreover, there are efficient algorithms [Sho90] to find such an irreducible polynomial.
@@
While our construction of Lemma˜5.1 follows the same outline of Theorem 5.12 in [GUV09], one subtle difference is that after every application of Lemma˜5.3 or ˜5.4, we replace the random source $X$ by its projection onto the dual space of the linear map.
@@
The difference between our construction and Theorem 5.12 in [GUV09] are (1) we replace every operation of leftover hash lemma by ˜5.4 and replace every operation of lossless condensers by Lemma˜5.3; (2) after each operation, we project the random source to the dual space of that operation. Specifically, let $X_{0}\in\mathbf{F}_{2}^{n}$ be the initial random source and $X_{i}\in\mathbf{F}_{2}^{n_{i}}$ be the random source after applying $i$ times ˜5.4 and Lemma˜5.3. For example, suppose the $(i+1)$-th operation applies the lossless condenser $h_{i}:\mathbf{F}_{2}^{n_{i}}\times\mathbf{F}_{2}^{d_{i}}\rightarrow\mathbf{F}_{2}^{m_{i}}$ with seed $s_{i}$ in Lemma˜5.3, say $h_{i}(X_{i},s_{i}):=A\cdot X_{i}$ for some full rank matrix $A\in\mathbf{F}_{2}^{m_{i}\times n_{i}}$. Then we set the next random source to be $X_{i+1}:=A^{\bot}\cdot X_{i}$ where $A^{\bot}\in\mathbf{F}_{2}^{(n_{i}-m_{i})\times n_{i}}$ is the dual of $A$. This works because Lemma˜5.3 and ˜5.4 hold for any $n$ and any $k<n$.
@@
The analysis of the error follows the same argument of [GUV09]. A small difference is to bound the min-entropy $X_{i+1}$ given $h_{i}(X_{i},s_{i})$. Since $h_{i}(\cdot,s_{i})$ is linear, $X_{i+1}$ is the exact distribution of $X_{i}$ conditioned on $h_{i}(X_{i},s_{i})$ such that $H_{\infty}\big(X_{i+1}\big)=H_{\infty}\big(X_{i}\mid h_{i}(X_{i},s_{i})\big)$. ∎
@@ Appendix A Connection between Min-wise Hash and PRG for Combinatorial Rectangles
Here we describe the connection between min-wish hash families and pseudo-random generators for combinatorial rectangles [SSZZ00, GY20].
@@
Plugging Theorem˜2.7 to Lemma˜A.1, Gopolan and Yehudayoff [GY20] had the following results of min-wise and $k$-min-wise hash with small errors.
@@
$=$ $\frac{1}{M}\sum_{\frac{|X|}{\ell}\cdot\frac{\theta}{M}\leq\ell^{-0.1}}(1-\theta/M)^{|X|-1}\cdot\left(1\pm 2^{-0.5C_{s}\cdot t}\right).$
@@
$\dfrac{1}{|X|^{3C-1}}+2^{-0.5C_{s}\cdot t}+2|X|\cdot\left(\exp(-\ell^{0.8})+N^{-O(1)}\right)<2^{-2C\cdot t}.$

### v3

@@ 2 Preliminaries
Explicit constructions of $t$-wise independent family of seed length $O(t\log NM)$ are well known. We state two useful bounds on $t$-wise independent random variables [minwise_Indyk, minwise_FPS11]. In the rest of this work, we use $\operatorname*{\mathbb{E}}_{X:t\text{-wise}}[f(X)]$ denote the expectation of $f(X)$ when $X$ is $t$-wise independent.
@@
PRGs for combinatorial rectangles have been extensively studied in the past few decades [ASWZ96_PRG_rect_comb, LLSZ95, Lu02, GMRTV, GY15]. Our constructicons of min-wise hash are based on these PRGs and their techniques. We state the state-of-the-art here by Gopalan and Yehudayoff, i.e., Theorem 1.9 in [GY15].
@@ 3 Min-wise Hash of Polynomial Size
The term in (3.3) is the error introduced by $\mathsf{PRG}_{1}$, which is multiplicative with respect to $\Pr[\sigma(y)=\theta]$. So the last piece is to show the total error of (3.3) over $\theta$, $\sum_{\theta\in[M]}\Pr_{\sigma\sim\mathcal{G}}[\sigma(y)=\theta\wedge\min\sigma(B_{j})>\theta]\cdot 2^{-C_{s}\cdot t}$, is bounded by $o(\delta)/|X|$. The observation here is that $\sum_{\theta\in[M]}\Pr_{\sigma\sim\mathcal{G}}[\sigma(y)=\theta\wedge\min\sigma(B_{j})>\theta]$ is the exact probability of event $(\sigma(y)<\min\sigma(B_{j}))$ under $(0.1C_{s}+1)$-wise independence since $\mathcal{G}$ is $(0.1C_{s}+1)$-wise independent. Thus by the classical result of Indyk [minwise_Indyk], this part is bounded by $O(1)/|B_{j}|$.
@@
Theorem 3.5 (Theorem 1.1 in [minwise_Indyk]).
@@
    After a union bound, with probability $1-1/\ell^{3C}$, $|B_{i}|\leq\ell^{0.1}+\operatorname*{\mathbb{E}}[|B_{i}|]\leq 2\ell^{0,1}$ for all $i$.
@@ 3.3.1 The First Part
We show (3.5) via the Nisan-Zuckerman framework. One subtle thing is that the extractor error $\mathsf{ExtErr}=2^{-C_{s}\cdot t}$ in (3.5) is multiplicative with respect to $\Pr_{\sigma:\text{$(0.1C_{s}+1)$-wise}}[\sigma(y)=\theta]=1/M$, different from applying the PRG fooling combinatorial rectangles directly. This is shown by permuting the order of $s_{1},\ldots,s_{\ell}$ such that $z_{j}:=\mathsf{Ext}(w,s_{j})$ is the first input of the read-once branching program.
Given the allocation $g$, we consider a width-$2$ length-$\ell$ read-once branching program $P$ whose input alphabet is $\{0,1\}^{0.3C_{e}\cdot\log N}$ corresponding to $z_{1},\ldots,z_{\ell}$. Because of the definition $B_{i}:=\{x\in X\setminus y:g(x)=i\}$ and $h^{\prime}(i):=\sigma_{g(i)}(i)$ for $\sigma_{1}:=G_{z_{1}},\ldots,\sigma_{\ell}=G_{z_{\ell}}$, we fix $g$ and rewrite $(h^{\prime}(y)=\theta\wedge\min h^{\prime}(X\setminus y)>\theta)$ as the conjunction of $\ell$ events depending on $z_{1},\ldots,z_{\ell}$:
@@ 3.3.2 The Second Part
$(1-\theta/M)^{|B_{i}|}\pm\varepsilon=(1-\theta/M)^{|B_{i}|}\cdot(1\pm 2\cdot 2^{-0.5C_{s}\cdot t}),\ \forall i\in[\ell].$
@@
$\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)=(1\pm 4\cdot 2^{-(0.5C_{s}-1)\cdot t})\cdot\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}(1-\theta/M)^{|X|-1}.$
@@
$\frac{1}{M}(1-\theta/M)^{|X|-1}=\frac{1}{M}\sum_{(1-\theta/M)^{C_{g}}\leq 2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)\leq 2^{-0.5C_{s}\cdot t/C_{g}}.$
@@
$=$ $\frac{1}{M}\left(\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)+\sum_{(1-\theta/M)^{C_{g}}\leq 2^{-0.5C_{s}\cdot t}}\prod_{i\in[\ell]}\left((1-\theta/M)^{|B_{i}|}\pm\varepsilon\right)\right)\pm 1/\ell^{3C}$
$=$ $\frac{1\pm 4\cdot 2^{-(0.5C_{s}-1)\cdot t}}{M}\cdot\sum_{(1-\theta/M)^{C_{g}}>2^{-0.5C_{s}\cdot t}}(1-\theta/M)^{|X|-1}\pm 2\cdot 2^{-0.5C_{s}\cdot t/C_{g}}\pm 1/\ell^{3C}.$
@@
$4\cdot 2^{-(0.5C_{s}-1)\cdot t}+\dfrac{1}{\ell^{3C-0.9}}+|X|\cdot 2^{-0.5C_{s}\cdot t/C_{g}}\leq 2^{-(0.5C_{s}-1)\cdot t+2}+2^{-2.5C\cdot t}+2^{-(0.5C_{s}/C_{g}-1)\cdot t}<2^{-2C\cdot t}.$
@@
$=$ $\dfrac{1}{M}\sum_{\theta\leq 0.5M\cdot\ell^{-0.2}}(1-\theta/M)^{|X|-1}\cdot\left(1\pm 2^{-0.001C_{s}\cdot t}\right).$
@@
    $\leq$ $2^{2S-(C_{s}-1)\cdot t\cdot S}=N^{-O(1)},$
    where we plug the definition of $S=C^{\prime}\cdot\log\log N$ in the last step.
@@
$\dfrac{1}{\ell^{3C-1.1}}+2^{-0.001C_{s}\cdot t}+|X|\cdot\left(\exp(-\ell^{0.6})+N^{-O(1)}\right)<2^{-2C\cdot t}.$
@@
where the last line applies the fact that only those $s_{j}=\alpha$ which make $(\sigma_{j}(y)=\theta\wedge\min\sigma_{j}(B_{j})>\theta)$ true have contributions to the expectation.
@@ 4 $k$-min-wise Hash
We use the second approach outlined in Section˜1.2 to construct $k$-min-wise hash in this section. Recall the definition that $\mathcal{H}=\{h_{i}:[N]\to[M]\}$ is a $k$-min-wise hash family of error $\delta$, if for any $X\subseteq[N]$ and $Y\subseteq X$ of size at most $k$, $\Pr_{h\sim\mathcal{H}}[\max h(Y)<\min h(X\setminus Y)]=\frac{1\pm\delta}{{|X|\choose|Y|}}$. Without loss of generality, we assume $|Y|=k$ in this section.
@@
If we neglect the $2^{-C_{s}\cdot t\cdot k+1}$ factor and consider $\sum_{\theta\in[M]}\Pr_{\sigma:(C_{e}+1)\cdot k\text{-wise}}[\max\sigma(Y)=\theta\penalty 10000\ \wedge\penalty 10000\ \min\sigma(B_{J})>\theta]$, this is the exact probability of $B_{J}\cup Y$ satisfying the $k$-min-wise hash condition under $(C_{e}+1)\cdot k$-wise independence. Feigenblat, Porat and Shiftan [minwise_FPS11] have shown this bounded by $2k!/|B_{J}|^{k}$ when $C_{e}$ is large enough. We state their results for completeness.
Theorem 4.6 (Theorem 1.1 in [minwise_FPS11]).
@@ 5 Extractors
We show how to design explicit extractors for Lemma˜5.1 in the rest of this section. Our construction will be based on the linear form of the lossless condenser in [GUV, CI_GUV_linear].
@@
We strengthen the linear seeded condensers in [GUV, CI_GUV_linear] such that they are surjective for every seed. We state the main properties of lossless condensers as follows, which reformulates Corollary 38 of [CI_GUV_linear] with an extra surjective guarantee.
@@
Let $\alpha>0$ be any constant. For any $n$, any $k<n$, and $\varepsilon>0$, there exists an explicit $(k,\varepsilon)$-lossless condenser $h:\mathbf{F}_{2}^{n}\times\mathbf{F}_{2}^{d}\to\mathbf{F}_{2}^{m}$ with $d\leq(1+1/\alpha)\cdot\log(nk/\varepsilon)+O(1)$ and $m\leq d+(1+\alpha)k$. Moreover, this lossless condenser satisfies the following two properties:
@@
Corollary 38 in [CI_GUV_linear] provides a linear lossless condenser with the above parameters although it is not necessarily surjective. In particular, its construction provides a condenser for every $n$ and every $k<n$ as long as there exists an irreducible univariate polynomial of degree $n$ in some finite extension field of $\mathbf{F}_{2}$. Irreducible polynomials of degree $n$ always exist because the number of irreducible polynomials of degree $n$ in the finite field $\mathbf{F}_{q}$ is $\frac{1}{n}\sum_{d\mid n}\mu(n/d)q^{d}$ greater than 0 (by the Gauss formula). Moreover, there are efficient algorithms [Shoup_irreducible] to find such an irreducible polynomial.
@@
While our construction of Lemma˜5.1 follows the same outline of Theorem 5.12 in [GUV], one subtle difference is that after every application of Lemma˜5.3 or ˜5.4, we replace the random source $X$ by its projection onto the dual space of the linear map.
@@
The difference between our construction and Theorem 5.12 in [GUV] are (1) we replace every operation of leftover hash lemma by ˜5.4 and replace every operation of lossless condensers by Lemma˜5.3; (2) after each operation, we project the random source to the dual space of that operation. Specifically, let $X_{0}\in\mathbf{F}_{2}^{n}$ be the initial random source and $X_{i}\in\mathbf{F}_{2}^{n_{i}}$ be the random source after applying $i$ times ˜5.4 and Lemma˜5.3. For example, suppose the $(i+1)$-th operation applies the lossless condenser $h_{i}:\mathbf{F}_{2}^{n_{i}}\times\mathbf{F}_{2}^{d_{i}}\rightarrow\mathbf{F}_{2}^{m_{i}}$ with seed $s_{i}$ in Lemma˜5.3, say $h_{i}(X_{i},s_{i}):=A\cdot X_{i}$ for some full rank matrix $A\in\mathbf{F}_{2}^{m_{i}\times n_{i}}$. Then we set the next random source to be $X_{i+1}:=A^{\bot}\cdot X_{i}$ where $A^{\bot}\in\mathbf{F}_{2}^{(n_{i}-m_{i})\times n_{i}}$ is the dual of $A$. This works because Lemma˜5.3 and ˜5.4 hold for any $n$ and any $k<n$.
@@
The analysis of the error follows the same argument of [GUV]. A small difference is to bound the min-entropy $X_{i+1}$ given $h_{i}(X_{i},s_{i})$. Since $h_{i}(\cdot,s_{i})$ is linear, $X_{i+1}$ is the exact distribution of $X_{i}$ conditioned on $h_{i}(X_{i},s_{i})$ such that $H_{\infty}\big(X_{i+1}\big)=H_{\infty}\big(X_{i}\mid h_{i}(X_{i},s_{i})\big)$. ∎
Acknowledgements
We appreciate anonymous reviewers for their helpful comments. We are also grateful to Parikshit Gopalan for pointing out some typos in an earlier version of this paper.
@@ Appendix A Connection between Min-wise Hash and PRG for Combinatorial Rectangles
Here we describe the connection between min-wish hash families and pseudo-random generators for combinatorial rectangles [SSZZ00, GY15].
@@
Plugging Theorem˜2.7 to Lemma˜A.1, Gopalan and Yehudayoff [GY15] had the following results of min-wise and $k$-min-wise hash with small errors.
@@
$=$ $\frac{1}{M}\sum_{\frac{|X|}{\ell}\cdot\frac{\theta}{M}\leq\ell^{-0.1}}(1-\theta/M)^{|X|-1}\cdot\left(1\pm 2^{-0.001C_{s}\cdot t}\right).$
@@
$\dfrac{1}{|X|^{3C-1}}+2^{-0.001C_{s}\cdot t}+|X|\cdot\left(\exp(-\ell^{0.8})+N^{-O(1)}\right)<2^{-2C\cdot t}.$